### v2

# Flow-Inspired Multi-Robot Real-Time Scheduling Planner

Abstract
Collision avoidance and trajectory planning are crucial in multi-robot systems, particularly in environments with numerous obstacles. Although extensive research has been conducted in this field, the challenge of rapid traversal through such environments has not been fully addressed. This paper addresses this problem by proposing a novel real-time scheduling scheme designed to optimize the passage of multi-robot systems through complex, obstacle-rich maps. Inspired from network flow optimization, our scheme decomposes the environment into a network structure, enabling the efficient allocation of robots to paths based on real-time congestion data. The proposed scheduling planner operates on top of existing collision avoidance algorithms, focusing on minimizing traversal time by balancing robot detours and waiting times. Our simulation results demonstrate the efficiency of the proposed scheme. Additionally, we validated its effectiveness through real world flight tests using ten drones. This work contributes an effective scheduling planner capable of meeting the real-time demands of multi-robot systems in obstacle-rich environments.
Code: https://github.com/chengji253/FRSP
I INTRODUCTION
Swarm robotics holds significant potential for various real-world applications, including air traffic control, search and rescue missions, and target detection. Much research has focused on ensuring the smooth and safe navigation of swarm robots in obstacle-rich environments.
In obstacle-rich environments, the primary challenge for multi-robot navigation is effective trajectory planning. Methods such as the velocity obstacle [1] [2], distributed model predictive control [3] [4], gradient-based local planning methods [5] and optimization-based methods [6] [7] have been proposed to address it. These approaches typically employ a two-stage process: front-end path planning for discrete path search and back-end collision avoidance optimization to refine trajectories. Common front-end path planning methods include point-to-point [1, 2, 3, 4], Astar [5], and methods like ECBS (Enhanced Conflict Based Search) [6] [7]. These studies mainly focus on the back-end, dealing with collision avoidance, reducing computational load, and minimizing deadlock. However, we find that there is still much room for improvement in the front-end path planning, especially in the case where multiple robots need to rapidly traverse an environment with obstacles.
Achieving rapid passage of multi-robot systems through obstacle-rich areas requires maximizing the utilization of every available path. This concept is similar to network flow problems, which aim to maximize flow from source to sink in a network [8]. The core of network flow lies in optimizing network capacity utilization to enhance flow and speed. We observe that the idea from maximum flow problems can be applied to efficiently schedule multi-robot systems, accelerating their traversal through obstacle-rich regions.
However, real-time and efficient scheduling of multi-robot systems in obstacle-rich environments poses significant challenges. Flow-based methods cannot be directly applied to multi-robot systems for several reasons. First, flow-based methods require a predefined network structure, but real-world map environments often lack such readily usable network structures. Second, these methods typically have long computation times and are often designed for offline use, lacking real-time applicability. Additionally, flow-based methods do not account for robot collision avoidance and motion models. In scenarios with a large number of robots, collision avoidance becomes a primary cause of congestion. At this point, robots face a decision: either wait in congested areas or choose a longer but uncongested path. Therefore, a new scheduling scheme is required, one that can determine in real time which robots should wait and which should reroute, while also specifying alternative routes.
To address these challenges, we propose a real-time scheduling scheme for multi-robot systems. Our planner builds on existing collision avoidance algorithms. It does not directly solve robot collision avoidance but performs real-time scheduling based on the map environment and robot status. The scheduling scheme first decomposes the initial map and extracts information to form a network containing nodes, edges, and capacity information. Subsequently, the planner calculates the path for each robot in real-time based on the current position of the robots and the congestion level of the map channels. The planner optimizes path selection from the perspective of the entire swarm, balancing detouring and waiting to minimize the time required for robots to traverse obstacle-rich environments. Experimental results show that our planner has high computational efficiency. And we conducted flight tests to demonstrate that the proposed scheme can be applied to real-world settings.
Our contributions can be summarized as follows:
  • We propose a scheduling planner that utilizes the traversable areas of the map, thereby reducing the traversal time for multi-robots in obstacle-rich environments.
  • Our scheduling planner exhibits high computational efficiency, meeting the real-time requirements of the robots. And we have validated the effectiveness of our algorithm through extensive experiments.
  • To the best of our knowledge, we are the first to apply the flow-based idea to multi-robot real-time scheduling, in order to improve the efficiency of traversing environments with multiple obstacles.
II related work
Numerous methods have been devised to tackle the navigation challenges of multiple robots in environments filled with obstacles, each with its own emphasis. Methods based on Velocity Obstacle (VO) and their derivatives have exhibited good performance in multi-robot navigation scenarios [1][2]. Luis et al.[3] introduced a distributed model predictive control algorithm for real-time trajectory generation of multiple robots. Soria et al.[4] presented a predictive model that optimizes potential field principles to improve the speed, order, and safety of aerial swarms in complex environments. The front-end paths in these studies [1, 2, 3, 4] are rather straightforward, typically being point-to-point paths. Zhou et al.[5] proposed a gradient-based local planning method for multi-robot systems, where the front-end path search uses Astar. Park et al.[6] proposed a method that combines ECBS and bernstein polynomial-based optimization for planning. Li et al.[7] proposed a planning method that merges ECBS with a priority-based nonlinear optimization approach. Both [6] and [7] use ECBS as a multi-agent path finding method for front-end search. Although these studies mainly concentrate on collision avoidance, reducing computational load, and handling complex scenarios, they overlook the time delays caused by congestion as the number of robots increases. As a result, when the map becomes more intricate and the number of robots grows, the efficiency of these front-end search methods drops significantly.
Since Ford and Fulkerson’s pioneering work on the maximum flow problem in networks [8], this area has received substantial attention. Variations like the multi-commodity flow problem share similarities with multi-robot navigation challenges. The multi-commodity flow problem aims to transport a set of commodities from sources to destinations while adhering to network capacity constraints and minimizing the overall transportation cost or time [9, 10, 11, 12]. However, these methods generally require prior knowledge of network information and focus on strictly meeting network capacity and boundary constraints. Their computational time is often long, rendering them unsuitable for direct application in multi-robot navigation. There are also some studies attempting to apply flow-based concepts to robot path planning. Yu et al.[13] showed that multi-agent path planning can be reduced to a network flow problem and proposed an algorithm with a complexity of $O(nVE)$. Dugas et al.[14] modeled crowds as a flow model and proposed a method for robot navigation within crowds. Janchiv et al.[15] applied flow network techniques to the multi-robot complete coverage path planning problem. Nevertheless, no research has yet applied flow-based ideas to multi-robot real-time scheduling and planning.
III Problem Statement
We consider a 2-dimensional workspace $\mathcal{W}\subseteq\mathbb{R}^{2}$ and a set of $M$ static obstacles $\mathcal{O}=\left\{o_{1},\ldots,o_{M}\right\}$ with $o_{i}\subset\mathcal{W}$. There are $N$ robots indexed by $n\in\{1,2,\ldots,N\}=\mathcal{N}$ who need to pass through the environment. Each robot $n$ has a start location $\mathbf{s}^{(n)}\in\mathcal{W}$ and a goal location $\mathbf{g}^{(n)}\in\mathcal{W}$. The robots are modeled as unit masses in $\mathbb{R}^{2}$, with single integrator dynamics. We use $\mathbf{p}^{(n)}(t),\mathbf{v}^{(n)}(t)$ to represent the discretized position, velocity at time step $t$ of the $n$th robot. With a discretization step $h$, the dynamic equations are given by
$\mathbf{p}^{(n)}(t+h)=\mathbf{p}^{(n)}(t)+h\mathbf{v}^{(n)}(t)$ (1)
We constrain the motion of the robots to match the physics of the vehicle. The robots have limited actuation, which bounds its maximum velocity: $\left\|\mathbf{v}^{(n)}(t)\right\|\leq\mathbf{v}_{\max}.$ The collision avoidance constraint is designed such that the robots safely traverse the environment:
$\left\|\mathbf{p}^{(i)}(t)-\mathbf{p}^{(j)}(t)\right\|\geq r_{\min},\;i,j\in\mathcal{N},i\neq j$ (2)
$\left\|\mathbf{p}^{(i)}(t)-o_{j}\right\|\geq r_{\min},\;i\in\mathcal{N},o_{j}\in\mathcal{O}.$ (3)
The Equs.(2), (3) show that the distance between any two robots and the distance between robots and obstacles must be higher than the minimum safety distance $r_{\min}$.
To fulfill the mission requirements, robots must navigate from starting points, traverse an obstacle-rich environment, and ultimately reach their destinations. During this process, robots must account for collision avoidance both among themselves and with the obstacles. The efficiency of the overall traversal of multi-robot system is measured by the time taken for the last robot to reach the destination.
Figure 1: A motivation example.
Fig.1 shows a motivation example. Ten robots need to move from the starting position to the target position. There are two selectable paths in the figure, namely Path1 and Path2. Path1 can only allow one robot to pass through each time, and it takes one second to pass. Path2 can allow two robots to pass through simultaneously each time, and it takes 3 seconds to pass. If all ten robots choose the shorter path Path1, then it will take ten seconds for all robots to reach the target position. However, if six robots are assigned to Path1 and four robots to Path2, then the six robots choosing Path1 will take six seconds to pass, and the four robots passing through Path2 simultaneously will also take six seconds. Thus, under this strategy, it only takes six seconds for all ten robots to pass through this area.
This example shows that by allocating some robots to paths with longer lengths but higher capacities, the overall efficiency can be improved. For the more complex situation (a more complex map, a larger number of robots, and the need to consider the motion model and collision avoidance). How to achieve real-time scheduling and planning, and reasonably allocate the number of robots according to the capacity of the paths to improve efficiency is not obvious. Therefore, a sophisticated scheduling scheme is needed.
IV Network construction
In this section, we will thoroughly explore the details of network construction. The initial map is typically a grid map that includes information about obstacles and passable areas. The network construction process involves extracting the passable area information from the grid map and then constructing a network with path and capacity information, which will serve as the foundation for subsequent scheduling. Note that unlike the planner that runs in real-time, the process of network construction only needs to be initialized once.
To build the subsequent network, we first employ the Boustrophedon Cellular Decomposition method [16] for map decomposition. This method scans the map in a back and forth motion, much like how oxen plow fields, dividing the map into multiple cells. Let the set of all cells be denoted as $\mathcal{C}$. Through map decomposition, the entire map is partitioned into multiple distinct regions, which we term cells. Each cell is connected to others and has adjacent boundary information. After processing this boundary information, we can introduce the concepts of $\operatorname{PathNode}$ and $\operatorname{PathPos}$.
Figure 2: The network construction process. The blue circles represent the $\operatorname{PathNode}$ of the network, and the red circles represent the $\operatorname{PathPos}$ of each node.
IV-A $\operatorname{PathNode}$ and $\operatorname{PathPos}$
$\operatorname{PathNode}$ and $\operatorname{PathPos}$ refer to some key positions on the adjacent boundaries of cells, which characterize the traffic capacity of the boundaries. First, we extract the length and position information of the boundaries of each cell, and determine the traffic positions on the common boundaries of adjacent cells. Denote the length of the common boundary as $L_{B}$. Then, the number of traffic positions that can be determined on this boundary is $\frac{L_{B}}{\alpha\;r_{\min}}$, where $\alpha$ is a constant, usually taking values between 1 and 2 to ensure that the robot has sufficient maneuvering space. The number of passage positions reflects the capacity of the boundary, that is, the maximum number of robots that can pass through simultaneously. These passage positions are defined as $\operatorname{PathPos}$ (shown as red circles in Figure (2)). Next, we group the passage positions on the same boundary. Every $N_{B}$ passage positions are grouped into one group, where $N_{B}$ represents the number of positions in each group. Subsequently, a new node is placed at the average position of each group, and this new node is defined as $\operatorname{PathNode}$ (shown as blue circles in Figure (2)). If the number of remaining red positions is less than $N_{B}$, these positions form a separate group. $\operatorname{PathNode}$ and $\operatorname{PathPos}$ play an important role as the bridge between the map and the network.
After establishing the nodes, we create links between the nodes on the upper and lower boundaries of a cell. The nodes on the lower boundary of a cell are connected to all nodes on the upper boundary, ultimately forming a complete network. Define the capacity of node $i$ as $\operatorname{Cap}(i)$, which is the number of passage positions of this node. Denote the network as a graph $\mathcal{G}=\{\mathcal{V},\mathcal{L}\}$ that includes a set of nodes $\mathcal{V}$ and a set of links $\mathcal{L}$. Each link $l=(i,j)\in\mathcal{L}$ has its corresponding length $\operatorname{Len}(l)$. The start of link $\operatorname{START}(l)=i$, the end node of link $\operatorname{END}(l)=j$. Let $\operatorname{UP}(c)\subset\mathcal{V}$ represent the upper boundary node set of cell $c$, $\operatorname{DN}(c)\subset\mathcal{V}$ represent the lower boundary node set of cell $c$. Define a function $\text{Dijkstra}(i,j)$ to find the dijkstra shortest path with a start node of $i$ and an end node of $j$. It will return a path node list connecting nodes $i$ and $j$. Denote the start cell and goal cell of robot $n$ as $c_{star}^{(n)},c_{goal}^{(n)}$.
Figure 3: The capacity of link
Define the capacity of link $l$ as $\operatorname{Cap}(l)$:
$\operatorname{Cap}(l)=\frac{\operatorname{Len}(l)\sin(\theta)}{\alpha\;r_{\min}}\min(\operatorname{Cap}(i),\operatorname{Cap}(j))$ (4)
where $i=\operatorname{START}(l),j=\operatorname{END}(l)$. $\theta$ is the angle between the link $l$ and the horizontal direction (Fig.(3)). The capacity of $l$ is the minimum capacity of its two endpoints, multiplied by the projection of the length of $l$ in the vertical direction, and then divided by $\alpha\;r_{\min}$. $\operatorname{Cap}(l)$ can be interpreted as the maximum number of robots that can be accommodated on link $l$ at one time.
IV-B State variable
After constructing the network, the next step is to define the state variables of the robots within the network. Let $c^{(n)}_{now}(t)$, $d^{(n)}_{pre}(t)$ and $d^{(n)}_{nex}(t)$ denote the current cell, previous node and target node of robot $n$ at time $t$. The link connecting the previous node and the target node is denoted as $l^{(n)}(t)$, with $\operatorname{START}(l^{(n)}(t))=d^{(n)}_{pre}(t)$ and $\operatorname{END}(l^{(n)}(t))=d^{(n)}_{nex}(t)$. The $l^{(n)}(t)$ represents the link where the robot is currently at time $t$. Denote the total number of robots on link $l$ as $\operatorname{Num}(l)$. Then, the state variable of the robot can be represented as:
$X^{(n)}(t)=\left(c^{(n)}_{now}(t),d^{(n)}_{pre}(t),d^{(n)}_{nex}(t),l^{(n)}(t)\right)$ (5)
where $c^{(n)}_{now}(t)\in\mathcal{C},d^{(n)}_{pre}(t)\in\mathcal{V},d^{(n)}_{nex}(t)\in\mathcal{V},l^{(n)}(t)\in\mathcal{L}$.
IV-C Control variable
In the network, the control variables for the robot $n$ at time $t$ are the path node list $\operatorname{PathNode}^{(n)}(t)$ and the path position list $\operatorname{PathPos}^{(n)}(t)$. The $\operatorname{PathNode}^{(n)}(t)$ represents the nodes that the robot needs to pass through, while $\operatorname{PathPos}^{(n)}(t)$ specifies the exact passage positions. During its movement, the robot progresses towards the first position in $\operatorname{PathPos}^{(n)}(t)$. Upon reaching the vicinity of this position, the following updates are triggered: (1) The first elements of the lists $\operatorname{PathNode}^{(n)}(t)$ and $\operatorname{PathPos}^{(n)}(t)$ are removed, indicating that the corresponding node and position have been traversed. (2) The previous node $d^{(n)}_{pre}(t)$ is updated to the node just traversed, and the next target node $d^{(n)}_{nex}(t)$ is updated to the first node in $\operatorname{PathNode}^{(n)}(t)$. (3) The robot selects the first position in the path position list $\operatorname{PathPos}^{(n)}(t)$ as the new target position and moves towards it. (4) As a result of the updates to the states $d^{(n)}_{pre}(t)$ and $d^{(n)}_{nex}(t)$, the edge $l^{(n)}(t)$ associated with the robot is also updated accordingly. Additionally, as the robot moves, it refreshes the state of the cell it occupies, and upon entering a new cell, the state $c^{(n)}_{now}(t)$ is automatically refreshed.
Figure 4: State and control variables example
Fig.(4) illustrates an example of a robot navigating within a network. The network consists of four nodes (blue) and their corresponding positions (red). The current state variable of the robot is $X^{(n)}(t)=\left(\text{Cell}_{1},V_{1},V_{2},L_{1}\right)$, indicating that the robot is currently at $\text{Cell}_{1}$, the previous node is $V_{1}$, the goal node is $V_{2}$, and the link is $L_{1}$. At this point, the path nodes $\operatorname{PathNode}^{(n)}(t)=[V_{2},V_{3},V_{4}]$, the path positions $\operatorname{PathPos}^{(n)}(t)=[P_{23},P_{33},P_{42}]$, and $\operatorname{Num}(L1)=1$. Upon reaching position $P_{23}$, the state of robot and control variables are updated. The new state variable becomes $X^{(n)}(t)=\left(\text{Cell}_{2},V_{2},V_{3},L_{2}\right)$. Subsequently, the path nodes $\operatorname{PathNode}^{(n)}(t)=[V_{3},V_{4}]$, the path positions $\operatorname{PathPos}^{(n)}(t)=[P_{33},P_{42}]$, and $\operatorname{Num}(L_{2})=1$.
V Scheduling Planner
In this section, we discuss the details of our scheduling planner. As illustrated in Fig.(5), our scheduling planner consists of three main components: path set search, path node selection and local position allocation. First, path set search will search for potential paths based on the current state of the robots in the network, and place these screened out paths in an alternative set. Subsequently, path node selection will comprehensively consider the path alternative sets of all robots, taking into account factors such as path length, path capacity, and the degree of robot congestion, and select appropriate path nodes for each robot. Finally, local position allocation will select a precise path position for each robot, that is, the passage position on the boundary. Through the collaborative operation of these parts, the control variables $\operatorname{PathNode}^{(n)}(t)$ and $\operatorname{PathPos}^{(n)}(t)$ are transmitted to the swarm robots, guiding them along their designated paths.
During movement, the robots continuously interact with the environment and with neighboring robots. In the presence of collision risks, the robots prioritize collision avoidance while following their paths. Both the scheduling planner and the collision avoidance algorithm operate in real-time, whereas the network construction is only required during the initial setup. Usually the frequency of the collision avoidance algorithm should be higher than the scheduling frequency.
Figure 5: The framework of our scheduling planner
V-A Path set search
In scheduling planner, to identify a path from the start point to the destination, we first need to generate a set of potential paths. Given that the number of paths increases exponentially with the scale of the network, directly searching all possible paths would render the problem intractable. Considering the real-time requirements of robotic systems, the number of paths included in the set must be limited; otherwise, the subsequent computation time would be excessively long. Therefore, our search algorithm must identify promising paths within a limited timeframe to facilitate subsequent optimization and selection.
Define the next-go-to path set for the $n$th robot as $\mathcal{P}^{(n)}_{\text{next}}(t)$, which represents the set of paths that the robot can potentially take from its current state in the future. During each decision-making instance, the robot considers only the paths within this set. The path search process is as follows: (1) Identify all $\operatorname{PathNode}$s on the upper boundary of the cell where the robot is currently located $\operatorname{UP}(c^{(n)}_{now}(t))$; (2) Identify all $\operatorname{PathNode}$s on the lower boundary of the goal cell $\operatorname{DN}(c_{goal}^{(n)})$; (3) Compute the paths from all upper boundary $\operatorname{PathNode}$s to the lower boundary $\operatorname{PathNode}$s using Dijkstra algorithm; (4) Add the resulting paths to path set. The next-go-to path set $\mathcal{P}^{(n)}_{\text{next}}(t)$ can be written as:
$\left\{\left[\text{Dijkstra}(i,j)\right]\mid i\in\operatorname{UP}(c^{(n)}_{now}(t)),j\in\operatorname{DN}(c_{goal}^{(n)})\right\}$ (6)
Thus, we find the next-go-to path set for the $n$th robot.
Intuitively, we can understand this search process in the following way. First, find out which exit nodes are available for selection in the cell where the current robot is located. Afterwards, determine which entrance nodes can be chosen to reach the target cell. Then, use the Dijkstra algorithm directly to find the shortest paths connecting the nodes, and add these paths to the set of selectable paths. We directly use the Dijkstra algorithm to find the shortest paths between nodes without considering the intermediate branching paths. The reason for this is that our algorithm is designed for real-time operation, rather than offline processing. Therefore, it is unnecessary to spend excessive time finding all possible intermediate paths. Instead, we only need to consider the viable paths from the robot’s current state to the destination. Our experiments have demonstrated that this search method is highly efficient, with the time for a single search involving 500 robots being approximately 0.4 seconds, and the final decision-making performance is satisfactory.
V-B Path node selection
After obtaining the result of $\mathcal{P}^{(n)}_{\text{next}}(t)$, the next step is to select a path from the set. The selected path result is assigned to $\operatorname{PathNode}^{(n)}(t)$ as the control variable of the robot. In this section, we will introduce a method based on mixed-integer programming.
Denote $p$ as a path element in the set $\mathcal{P}^{(n)}_{\text{next}}(t)$. Denote the decision variable as $z_{p}^{(n)}(t)$, which is a binary variable that represents whether the $n$th robot choose the path $p$ at time $t$. Then the set of decision variables $\mathcal{Z}$ is written as:
$\mathcal{Z}=\left\{z_{p}^{(n)}(t)\mid n\in\mathcal{N},p\in\mathcal{P}^{(n)}_{\text{next}}(t)\right\}$ (7)
The set $\mathcal{Z}$ has decision variables of all robots. Since the robot can only choose one path at a time, we need to introduce the constraint,
$\sum_{p\in\mathcal{P}^{(n)}_{\text{next}}(t)}z_{p}^{(n)}(t)=1.$ (8)
which means that the $n$th robot can only choose one path from the set $\mathcal{P}^{(n)}_{\text{next}}(t)$.
Denote the $\mathcal{P}_{\text{next}}(t)$ as the union of all sets $\mathcal{P}^{(n)}_{\text{next}}(t)$,
$\mathcal{P}_{\text{next}}(t)=\bigcup\left\{\mathcal{P}^{(n)}_{\text{next}}(t),n\in\mathcal{N}\right\}$ (9)
Denote the first and second link of path $p$ is $\operatorname{Fir}(p)$ and $\operatorname{Sec}(p)$. They are the link connecting the first node and the second node, and the link connecting the second node and the third node in the path $p$ respectively. Denote the $\operatorname{Fir}(\mathcal{P}_{\text{next}}(t))$ and $\operatorname{Sec}(\mathcal{P}_{\text{next}}(t))$ as the set of first links and second links of path set $\mathcal{P}_{\text{next}}(t)$,
$\operatorname{Fir}(\mathcal{P}_{\text{next}}(t))=\left\{\operatorname{Fir}(p)\mid p\in\mathcal{P}_{\text{next}}(t)\right\}\subset\mathcal{L},$ (10)
$\operatorname{Sec}(\mathcal{P}_{\text{next}}(t))=\left\{\operatorname{Sec}(p)\mid p\in\mathcal{P}_{\text{next}}(t)\right\}\subset\mathcal{L}.$ (11)
Denote $\mathcal{P}^{l}_{\text{fir}}(t)$ as the set of path that the first link of path $p$ is $l$,
$\mathcal{P}^{l}_{\text{fir}}(t)=\left\{p\mid p\in\mathcal{P}_{\text{next}}(t),\operatorname{Fir}(p)=l\right\}.$ (12)
Denote $\mathcal{P}^{l}_{\text{sec}}(t)$ as the set of path that the second link of path $p$ is $l$,
$\mathcal{P}^{l}_{\text{sec}}(t)=\left\{p\mid p\in\mathcal{P}_{\text{next}}(t),\operatorname{Sec}(p)=l\right\}.$ (13)
Introducing the definitions of these sets serves as the foundation for our subsequent introduction of queueing cost. We will analyze the potential for future congestion in robots based on whether all paths share the first and second links.
Figure 6: An example of first and second link. The blue circles represent the “$\operatorname{PathNode}$”. The red, orange and green lines represent three paths. The dotted line represents the link.
Fig.(6) shows an example of first link and second link. There are three paths $p_{1}=[V_{1},V_{2},V_{3},V_{4}]$ (red), $p_{2}=[V_{1},V_{2},V_{3},V_{6}]$ (orange), $p_{3}=[V_{5},V_{2},V_{3},V_{6}]$ (green). $p_{1},p_{2}$ share links $L_{1},L_{2}$. $p_{2},p_{3}$ share link $L_{2}$. At this time, $\mathcal{P}_{\text{next}}(t)=\left\{p_{1},p_{2},p_{3}\right\}$. The first links of all paths have $L_{1},L_{4}$, the second link of all paths is $L_{2}$. Thus, we have $\operatorname{Fir}(\mathcal{P}_{\text{next}}(t))=\left\{L_{1},L_{4}\right\}.$ $\operatorname{Sec}(\mathcal{P}_{\text{next}}(t))=\left\{L_{2}\right\}$. $\mathcal{P}^{l}_{\text{fir}}(t)=\left\{p_{1},p_{2}\right\},l=L_{1}$. $\mathcal{P}^{l}_{\text{fir}}(t)=\left\{p_{3}\right\},l=L_{4}.$ $\mathcal{P}^{l}_{\text{sec}}(t)=\left\{p_{1},p_{2},p_{3}\right\},l=L_{2}$.
V-B1 Link queuing cost
In this part, we will introduce how to calculate queueing cost through the first and second links. Define the queuing cost of the first link as $f_{\operatorname{Fir}}(t)$:
$\sum\limits_{l\in\operatorname{Fir}(\mathcal{P}_{\text{next}}(t))}\!\!\!\!\!\frac{\left(\sum\limits_{n\in\mathcal{N}}\sum\limits_{p\in\mathcal{P}^{l}_{\text{fir}}(t)}z_{p}^{(n)}(t)+\operatorname{Num}(l)-\operatorname{Cap}(l)\right)^{2}}{\operatorname{Cap}(l)^{2}}.$ (14)
The term $\sum_{n\in\mathcal{N}}\sum_{p\in\mathcal{P}^{l}_{\text{fir}}(t)}z_{p}^{(n)}(t)$ represents the number of times link $l$ is selected across all robots path results. When added to the current number of robots on link $l$, this gives the total number of robots on link $l$ after decision-making. We then subtract the capacity of link $l$, and compute the square of the resulting value. This represents the queuing cost for the first link. Summing the queuing cost for all first links yields $f_{\operatorname{Fir}}(t)$. This cost term ensures that the path node selection does not exceed the capacity of the links. When there are already many robots on an link, the remaining capacity is limited, and assigning more robots to that link will increase the loss. Similarly, if the sum of the assigned and existing robots is less than the capacity, the loss will also increase, encouraging the decision-making process to fully utilize the link capacities. Define the queuing cost of the second link as $f_{\operatorname{Sec}}(t)$:
$\sum\limits_{l\in\operatorname{Sec}(\mathcal{P}_{\text{next}}(t))}\!\!\!\!\!\!\frac{\left(\sum\limits_{n\in\mathcal{N}}\sum\limits_{p\in\mathcal{P}^{l}_{\text{sec}}(t)}z_{p}^{(n)}(t)+\operatorname{Num}(l)-\operatorname{Cap}(l)\right)^{2}}{\operatorname{Cap}(l)^{2}}.$ (15)
The Equ.(15) is identical to the first link cost of Equ.(14). By incorporating the cost of the second link, the length of the paths considered in each planning step is extended, thereby imparting a degree of predictiveness to the decision-making process. Typically, the coefficient of $f_{\operatorname{Sec}}(t)$ is smaller than $f_{\operatorname{Fir}}(t)$, as the congestion in the map changes with the movement of robot. Therefore, the cost of the first link is prioritized.
V-B2 Running cost
Running cost must also be considered. Each robot needs to balance between rerouting and waiting. Sometimes, the time cost of choosing a longer path may exceed the cost of waiting in place, necessitating a trade-off between the two. Therefore, a running cost penalty should be incorporated into the objective function to achieve more rational decision-making. Denote the running cost function as $f_{\operatorname{run}}(t)$:
$f_{\operatorname{run}}(t)=\sum_{n\in\mathcal{N}}\sum_{p\in\mathcal{P}^{(n)}_{\text{next}}(t)}\operatorname{Len}(p)z_{p}^{(n)}(t)$ (16)
where $\operatorname{Len}(p)$ is the length of path $p$. Putting these things together, a complete mixed-integer quadratic programming (MIQP) model is obtained, summarized below.
$\underset{\boldsymbol{z}_{p}^{(n)}(t)}{\operatorname{minimize}}\quad k_{1}f_{\operatorname{Fir}}(t)+k_{2}f_{\operatorname{Sec}}(t)+k_{3}f_{\operatorname{run}}(t)$ (17)
subject to: Equs.(6), (8), (9), (10), (11), (12), (13). $k_{1},k_{2},k_{3}$ are weight coefficients. After this MIQP process, the specific binary values $z_{p}^{(n)}(t)$ can be obtained, which represent the path node selection result $\operatorname{PathNode}^{(n)}(t)$.
V-C Local position allocation
Upon obtaining the path node selection results $\operatorname{PathNode}^{(n)}(t)$, the next step is to allocate specific node positions $\operatorname{PathPos}^{(n)}(t)$ for the robots. Each node has multiple positions, and we adopt a method of selecting the positions with the closest distance one by one. Initially, we select the position closest to the current position of robot. For each subsequent node, we select the position that is nearest to the previously allocated position. This process is repeated iteratively, ensuring each chosen position is the closest to the last assigned position. In this section, we directly choose the position based on the closest distance, rather than more complex methods. We found in the experiments that this assignment method works very well. When robots cannot reach the ideal positions due to congestion and choose the positions closest to themselves, the previous section “Path node selection” will replan new nodes for them, thus improving the overall efficiency. Therefore, this algorithm is simple and efficient.
VI Experiments
((a)) $t=1.4s$
((b)) $t=4.8s$
((c)) $t=6.2s$
((d)) $t=10.4s$
Figure 7: The scheduling process of 500 robots in the forest map.
In this section, we conduct two sets of experiments to explore the performance of our method in multi-robot planning. All experiments are carried out on a PC equipped with an Intel Core-i7 12700 CPU and 32GB of RAM.
The number of robots in the first set of experiments ranges from 10 to 100. (10, 20, 30, 40, 50, 60, 70, 80, 90, 100) For each number of robots, we conduct experiments on two different map types: random forest and maze. Five different instances are generated for each map type. In this set of experiments, the compared algorithms are as follows:
  • Astar: Each robot uses the Astar algorithm to calculate the path from the starting point to the goal point.
  • ECBS (Enhanced Conflict Based Search) [17]
  • P-SIPP (Prioritized Safe Interval Path Planning) [18]
We use FRSP to denote our algorithm. The coefficients of FRSP are set to $k_{1}=1,k_{2}=0.5,k_{3}=0.5$. The MIQP is solved by Gurobi [19].
In the second set of experiments, we explore the impact of a large number of robots on the algorithms. The number of robots is set to 100, 150, 200, 250, 300, 350, 400, 450, and 500. We also conduct experiments on two types of maps: forest and maze. However, during the experiment, it is found that when the number of robots is higher than 100, the success rates of the ECBS and P-SIPP algorithms are extremely low, making effective comparison impossible. Therefore, in this set of experiments, the following three methods are selected for comparison:
  • Astar: Each robot uses the Astar algorithm to calculate the path from the starting point to the goal point.
  • Greedy: Find all paths from the upper boundary node of the starting cell to the lower boundary node of the goal cell. Then, use a greedy approach to evenly distribute robots across each path according to the path length and capacity.
  • RunCost: The method is almost the same as FRSP. But the objective function only has running cost. The coefficients are set to $k_{1}=0,k_{2}=0,k_{3}=1$.
Two sets of experiments are set up with the aim of demonstrating that our algorithm not only outperforms the commonly used ECBS and P-SIPP algorithms within the number range of 10 to 100. In the large scale experiments (100 to 500) where both ECBS and P-SIPP are unable to cope, our algorithm can still handle and exhibit good performance.
The parameters for the experiments are set as follows: $r_{\min}=0.4$m, $\alpha=2$, $h=0.01s$, $\mathbf{v}_{\max}=3$m/s, $N_{B}=4$. The initial and goal positions of the robots are uniformly arranged in a matrix queue format. Our local collision avoidance algorithm utilizes the Reciprocal Velocity Obstacles (RVO) [1] method, chosen for two primary reasons. Firstly, RVO demonstrates reliable performance in multi-robot collision avoidance, reducing the likelihood of deadlocks. Secondly, it offers high computational efficiency, maintaining acceptable computation times even with a large number of robots. The frequency of collision avoidance algorithm is 100 Hz and the scheduling is 1 Hz.
((a)) Forest maps
((b)) Maze maps
Figure 8: Time results w.r.t. number of robots (10 to 100)
((a)) Forest maps
((b)) Maze maps
Figure 9: Time results w.r.t. number of robots (100 to 500)
Fig.8 shows the time results for the number of robots ranging from 10 to 100 in the forest and maze maps, while Fig.9 shows the time results for the number of robots ranging from 100 to 500 in the forest and maze maps. For every number of robots $N$, mean values are reported for each type of 5 maps. We can observe that: (1) The time result of FRSP is the best, while that of P-SIPP is the worst. The time results of ECBS and Astar are similar, with ECBS being slightly inferior to Astar. This indicates that the FRSP can better handle the congestion caused by the increase in the number of robots. (2) The Astar algorithm searches for the shortest path to the destination for each robot. ECBS and P-SIPP, on the other hand, search for discrete, conflict-free paths for the robots. However, these paths do not take into account the impact of congestion. As a result, as the value of $N$ increases, the time efficiency decreases. (3) FRSP still performs well in the range of 100 to 500. The performance of Greedy algorithm is the worst. (4) When $N$ is small, there is not much difference between the time of four methods. This is because when the number is small, congestion rarely occurs, so the time results of the methods are similar. (5) As $N$ increase, the time results increase. The gap between time results of FRSP and other methods is gradually increasing. This is because as $N$ increases, the congestion starts to worsen, and the robots will waste more time waiting. FRSP can plan new paths for the waiting robots, thus reducing the time it takes to cross. The larger $N$ is, the better FRSP works. (6) FRSP performs better on forest maps than on maze maps. This is because the forest map is more complex than the maze map, and there are more forks in the network structure. This shows that FRSP can adapt to complex map structures, and the more complex the map structure, the better the effect.
The average improvement
Range from 10 to 100 Range from 100 to 500
Astar ECBS P-SIPP Astar Greedy RunCost
FRSP Forest 9.04% 11.37% 14.78% 8.66% 7.52% 10.77%
Maze 7.07% 8.13% 11.57% 6.58% 4.68% 7.54%
TABLE I: The average improvement of FRSP
Average computation time (unit: seconds)
Number 10 50 100 200 300 400 500
Path set
search
0.050 0.076 0.105 0.158 0.262 0.326 0.456
Path node
selection
0.014 0.058 0.114 0.181 0.243 0.317 0.447
Position
allocation
0.000 0.002 0.004 0.008 0.015 0.013 0.025
Sum 0.064 0.136 0.223 0.347 0.520 0.656 0.928
TABLE II: The average computation time of FRSP
Table.I shows the average improvement of FRSP. It can be seen that: (1) the Greedy method is not as good as RunCost and Astar. This is because both RunCost and the Astar choose the shortest distance path, and they will wait when faced with congestion. But Greedy assigns robots to each path according to the greedy idea. Although congestion was reduced, time was wasted on further travel, suggesting that sometimes waiting may be a good option. (2) In the scenario of multi-robots moving from one side to the other side of an obstacle-filled area, the ECBS and P-SIPP algorithms perform even worse than the Astar algorithm. This is in line with expectations because in the algorithm design of ECBS and P-SIPP, the initial and goal positions of the robots are scattered, and these scattered positions are less likely to cause congestion. This indicates that the commonly used ECBS and P-SIPP are highly unsuitable for addressing congestion situations.
Table.II shows the computation time with respect to the number of robots $N$. It can be seen that: (1) Path set search and path node selection take up most of the computation time, while local position allocation takes very little computation time. (2) When $N$ is 10, the computation time of FRSP is about 0.06 seconds. When $N$ is 500, the computation time of FRSP is about 0.9 seconds. This shows that our FRSP is computationally efficient and meets the real-time requirements.
VI-A Real World Flight Test
Figure 10: Real flight test scene.
Figure 11: The flight process of 10 drones.
We conduct flight tests with ten Dij Tello drones within a forest environment. The drones are constrained to a space of $15m\times 5m\times 2m$ to prevent them from completely circumventing obstacle areas. We use OptiTrack to provide high-precision positioning information for drones. Fig.10 shows our real flight test scene where we placed 5 boxes as obstacles in the scene. The parameter settings of the real flight scene are as follows: $r_{\min}=0.5$m, $\alpha=1$, $h=0.01s$, $\mathbf{v}_{\max}=1$m/s, $N_{B}=1$. The frequency of collision avoidance algorithm is 100 Hz and the scheduling is 1 Hz. Fig.11 shows the flight process of 10 drones. All calculations are performed on a single laptop and no collisions or deadlocks occurred during the entire flight. The detailed process can be seen in the supplementary video.
VII Conclusion
In this paper, we proposed a flow inspired scheduling planner to optimize the traversal of multi-robot systems through obstacle-rich environments. Simulation and experimental results demonstrate the effectiveness of our scheduling planner. In the future work, we will consider how to perform map decomposition and network construction under obstacles of arbitrary shapes. Try different collision avoidance algorithms to enhance the applicability of our scheduling planner. And in the real world experiment, we will try to conduct more drones tests and try more complex environments.